one base of a trapezoid is three less than twice the length of the other base. The altitude of the trapezoid is 14 inches. if the area of the trapezoid is 63 square inches, find the length of each base
Answer by mestrydilip(12)

Length of one base = x
Lenght of other base = 2x - 3
Altitude = 14
Area = 1/2 (sum of the bases) X Altitude
63 = 1/2 ( x + 2x - 3) X 14
63 = 1/2 (3x - 3) X 14
63 = 7 (3x - 3)
9 = 3x - 3
12 = 3x
x = 4
other base = 2x - 3 = 2(4)-3 = 8 - 3 = 5